OpenStudy (anonymous):

Is there anyone who can help me with finding the slope of an equation? I will give medal.

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

i will

OpenStudy (anonymous):

me too

OpenStudy (alexandervonhumboldt2):

y=ax+b a is the slope

OpenStudy (anonymous):

To find the slope of an equation must use the formula y2-y1/x2-x1.

OpenStudy (alexandervonhumboldt2):

it is not on grath @SethGP

OpenStudy (alexandervonhumboldt2):

it is th eequation

OpenStudy (anonymous):

(-4,-1) and \[ 1\frac{ 1 }{ 2 } ,2\]

OpenStudy (alexandervonhumboldt2):

ok so yeah use slope fromua

OpenStudy (alexandervonhumboldt2):

(y1-y2)/(x1-x2)

OpenStudy (alexandervonhumboldt2):

(-1-2)/(-4-1.5)

OpenStudy (alexandervonhumboldt2):

-6/11

OpenStudy (anonymous):

so tht is the slope?

OpenStudy (alexandervonhumboldt2):

-6/11

OpenStudy (alexandervonhumboldt2):

nono 6/11

OpenStudy (alexandervonhumboldt2):

6/11

OpenStudy (alexandervonhumboldt2):

-/-=+

OpenStudy (alexandervonhumboldt2):

6/11

OpenStudy (anonymous):

ok thanks so -2-(-1)/-4-1.5 = 6/11

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

So I was right

OpenStudy (anonymous):

hey @SethGP do you know how to turn this into standard form? y-2=6/11(x-(-4))

OpenStudy (anonymous):

i can make this a new question so you can have a medal for helping me if u want?